ACCURATE DETERMINATION OF THE GROUND STATE CONSTANTS, INCLUDING A_{0} AND D_{K}^{0}, OF D_{3}Si^{35}Cl

A. CEAUSU, \underline{G. GRANER}, Laboratoire de Physique Moléculaire et Applications, CNRS, Bât. 350, Campus d’Orsay, F-91405 Orsay Cédex, France; E.B. MKADMI, and H. BÜRGER,, Anorganische Chemie, FB 9, Universität-Gesamthochschule, D-42097 Wuppertal, Germany.

In a previous study^{a}of monoisotopic D_{3}Si^{35}Cl, the analysis of the fundamental bands ν_{3} and ν_{6}, linked by a Coriolis resonance, was reported. New Fourier transform infrared spectra of this compound were recorded in the 500 and 1000 cm^{−1} regions. The first purpose was to obtain accurate ground state constants. The ’normal’ G.S. constants B_{0}, D_{J}^{0}, and D_{JK}^{0} were deduced from GSCD. As for the constants A_{0} and D_{K}^{0}, they were obtained by the now well-established method of combining data from the fundamental ν_{6}, the overtone (2ν_{6})^{∓2} and the hot band (2ν_{6})^{±2}-(ν_{6})^{±1}. Special care was taken to correct calibration differences between the two spectral regions. This was done by checking closed loops such as ν_{3}, 2ν_{3}, and 2ν_{3} - ν_{3} on one hand or ν_{6},(2ν_{6})^{0} and (2ν_{6})^{0} - ν_{6} on the other hand. The values obtained are A_{0} =1.427830 cm^{−1} and D_{K}^{0}=5.36× 10^{−6} cm^{−1} (provisional values). Some sextic H constants could also be determined. In a second stage, the upper state energies of (2ν_{6})^{±2}, (2ν_{6})^{0}, 2ν_{3} and ν_{3} + ν_{6} were fitted together to obtain the upper state constants. Note that, due to the weakness of the ν_{3} + ν_{6} band, the hot bands ν_{3} + ν_{6} - ν_{3} and ν_{3} + ν_{6} - ν_{6} were used instead.

aJ.E. Gadhi, H. Bürger, and E.B. Mkadmi, submitted to J. Mol. Spectrosc.